OpenStudy (anonymous):

on a number line max places a mark 3 units to the left of the point halfway between 1 and -3. on what point is maxs mark? 2 0 1 -1 -3 -4

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

well, what is half way between 1 and -3?

OpenStudy (anonymous):

i did the work yesterday and my answer was -4

OpenStudy (anonymous):

but i needed reassuance

OpenStudy (amistre64):

show me your work and ill reassure you :)

OpenStudy (amistre64):

in general, we find the midpoint and add the desired amount it moves: given an interval from a to b, the midpoint is: (a+b)/2 moving it by some amount, k, is therefore:\[\frac{a+b}{2}+k\] \[\frac{1+(-3)}{2}+(-3)\] \[\frac{-2}{2}-3\] \[-1-3\]

OpenStudy (anonymous):

im confused now all i did was go to a graph and find the point i n between 1 and -3 and then move 3units left

OpenStudy (amistre64):

lol, thats all i did too, but with algebra instead

OpenStudy (amistre64):

the midpoint is -1, and move 3 to the left ...

OpenStudy (anonymous):

so -4?